Carbon monoxide and chlorine gas react to form phosgene:
CO(g)+Cl2(g)?COCl2(g) Kp = 3.10 at700 K

If a reaction mixture initially contains 172torr of CO and 357torr of Cl2, what is the mole fraction of COCl2when equilibrium is reached?
Answer #1

CO (g) + Cl2(g) ------> COCl2(g)

K_{P}=\frac{P_{COCl_{2}}}{P_{CO}\times P_{Cl_{2}}}

Kp = 3.10

PCO = 172 torr = 170/760 atm = 0.224 atm

PCl2 = 357 torr = 375/760 atm =0.493 atm

3.10=\frac{P_{COCl_{2}}}{0.224atm\times 0.493atm}

PCOCl2 = 0.342 atm

Total pressure at equilibrium is the sum of partial pressures of the gases CO, Cl2 and COCl2.

Ptot = 0.224 + 0.493 + 0.342 = 1.059 atm

For COCl2, the partialpressure is the product of its mole fraction and total pressure.

PCOCl2= XCOCl2.Ptot

(0.342 atm) = XCOCl2(1.059 atm)

XCOCl2 = 0.322

Therefore, molefraction of COCl2 at equilibrium is 0.322.